# Designing and Evaluating AI Margin Notes in Document Reader Software

### v1

@@ 4. Experimental Method
We conducted a series of experiments to explore the design parameters of AI margin notes. Specifically, we were interested in understanding how integration, selection automation, and human and AI involvement affect reading comprehension and duration, psychological ownership, task workload, and user preferences. All experiments were conducted using the Prolific crowdsourcing platform. [^1] https://www.prolific.com
@@ 4.1. Task
The primary experimental task was composed of two stages. First, a reading stage, where participants read non-fiction documents that were approximately 500 words each. For each document, they interacted with an LLM assistant using a different technique that represents a variation within the design parameter under evaluation. Second, a test stage, where participants completed reading comprehension tests about each document. Each test consisted of six multiple choice questions. The documents and questions were developed by Wallace et al. (Wallace et al., 2022) and were designed by a learning and reading specialist to be suitable for an eighth grade reading level. [^2] Note that we slightly modified the wording of some questions, as we noticed that some questions could be successfully answered by looking at the wording of other questions. At the time of writing this paper, Wallace et al.’s license agreement permits reuse, modification, public display, and redistribution for non-commercial research purposes. This is representative of documents targeted for the general public (Nielsen, 2015).
@@ 4.2. Apparatus
Our experimental system was a custom Node.js and React web application that implemented two types of interfaces, one for each stage: the reading interface, where participants read a document and interacted with an LLM using a specific technique (Figure 5); and the testing interface, where participants answered time-bounded, multiple-choice reading comprehension tests.
@@ 4.2.1. Reading Interface
The document was displayed on the left side of the screen. The specific design variation of the experimental condition was displayed beside it (details provided in each individual experiment). All design variations that involved an LLM used GPT 4.1 mini. The top of the screen contained a toolbar indicating how many comments or responses they still had to complete and a blue “Finish Reading” button to end the trial.
@@ 4.2.2. Testing Interface
[…]
Interface with a header prompting users to create comments. On the left, text titled ”Flowering Plants” explains their structure. Right side has a blank space.
@@ 4.3. Procedure
Participants received a link to an experimental web application through Prolific. They had to use a desktop or laptop computer, which was strictly enforced through the web application. First, they entered basic demographic information and read instructions. Next, they completed the reading stage for all documents. There was no time limit for the reading stage, however, to focus on the effect of the different techniques rather than the number of responses, participants could only leave 3 AI margin notes (or receive 3 responses from the AI assistant, depending on the technique). After reading the document, they answered 10 questions about their experience and 1 question about their prior knowledge of the document’s content. They repeated this for the other documents, then answered 2 questions about their overall preferences. They also described other study aids they used for the test, such as taking a screenshot of the document or writing notes outside of the reading interface.
@@
Two hours later, the participant was invited to return for the test stage. Given the lower reading difficulty of the documents, we increased the difficulty of the test by making it closed-book and restricted to 60 seconds.
4.4. Design
All experiments use within-subjects design since the documents were relatively short, and we wanted participants to compare their experiences with each technique. There is one independent variable, condition, which was randomly assigned. For a single condition, one of six documents from Wallace et al. (Wallace et al., 2022) was randomly-assigned. For increased internal validity, a single document was not restricted to a single condition and could be assigned to any condition.
[…]
Two interfaces: (a) a chat-based interface with a question about flower roles, and (b) an AI margin note with the same content displayed in a single comment box.
4.5. Measures
Reading comprehension and duration were calculated from logs, and all other metrics were calculated from subjective questionnaires that participants completed after the reading stages. All used a 0-100 interval range.
4.5.1. Reading Comprehension
4.5.2. Duration
4.5.3. Psychological Ownership
This was measured by asking two quantitative questions about Personal Ownership and Responsibility (Caspi and Blau, 2011). The average of the two create a composite measure, as done in prior work (Joshi and Vogel, 2025a, b). The internal consistency reliability score, which describes how consistently different items on a questionnaire describe the same underlying concept, was high for all experiments (.80 $\leq\alpha\leq$ .90), suggesting that the composite measure was appropriate to use for data analysis. As described previously, this is an important measure since fostering psychological ownership can improve learner experiences.
4.5.4. Task Workload
We used factors from the NASA-TLX: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Techniques that are more manual or that require more human involvement may involve more mental demand and effort, which could be beneficial for learning (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a; Tyler et al., 1979), however, this may not happen if they are too frustrating (Grund et al., 2024).
4.5.5. Preferences
We asked about Frequency of Use, a question from the system usability scale (SUS) representing how frequently the participant would use the feature if made available to them. After trying all techniques, participants also gave each condition a Ranking, where 1 was the best, 2 was the second best, and so on. Ties were allowed. To establish an Overall Ranking of conditions, we use the Condorcet voting method (Young, 1988). An overall rank of 1 means that technique ‘defeats’ all others in pairwise comparisons. An overall rank of 2 means that technique ‘defeats’ all others except for the technique ranked 1, and so on. This is important to study, as users would likely want to keep using techniques that they like, which is especially useful if it also improves their learning outcomes.
5. Experiment 1: Integration
5.1. Participants
We recruited 29 participants through Prolific. Participants were restricted to the United States and Canada, and those who had completed 2,500 previous tasks on the platform with a 99-100% approval rating. Three participants (10%) were removed for using other study aids prior to the test, leaving 26 valid responses (Table A.1). All self-reported proficiency in reading in English. Participants received \$10 in total, and the top 25% scoring participants received an additional \$5 bonus to incentivize them to do well on the test. The experiment took roughly 25 minutes in total.
5.2. Apparatus
5.3. Results
We use Wilcoxon signed-rank tests to investigate the effects of condition on the various measures. To streamline the presentation of results, details of these statistical tests are shown in Table A.2.
5.3.1. Preferences
Although there were no differences for the aforementioned measures, most participants preferred prompting with AI margin notes. From the Condorcet voting method, we observed that note received an Overall Ranking of 1, and chat received an Overall Ranking of 2. Notably, majority of participants (17, 65%) assigned note a Ranking of 1, and chat a Ranking of 2 (16, 62%).
To better understand why participants preferred note over chat, we examined participants’ free-form responses. Several (12, 46%) mentioned how note was easier and more intuitive than chat. Participants described how they “liked [not] having to move to a separate chat”(P21), how “selecting text [was] more natural when one has a question related to that part of the text”(P14), and how “having the comment embedded right into the document [was] handy and easy to reference along with the actual text”(P5).
[…]
Two processes for generating AI margin notes. In (a), automatic text selection generates notes after a button press. In (b), manual text selection generates notes from highlighted segments. Both methods produce concise notes summarizing flower structures and functions.
Overall, our results suggested that the integration of AI margin notes is highly desirable and preferable over chat-based interfaces as it prevented switching between interfaces, it was easier for readers to indicate text they had questions about, and it was to reference individual comments later. In this experiment, associating text to the AI margin note was done manually by the participant. However, selecting text could also be done automatically by the LLM, which we explore further in the following experiment.
6. Experiment 2: Selection Automation
The goal of this experiment is to understand the effect of selection automation when creating AI margin notes. Participants either manually selected what text should be summarized (manual), or pressed a button to let the AI assistant place three summary AI margin notes in the document (automatic).
6.1. Participants
Using the same inclusion criteria as Experiment 1, we recruited 32 new participants on Prolific. Two participants (6%) were excluded for not attempting to answer any reading comprehension questions, or for using other study aids, leaving 30 valid responses (Table A.3). All self-reported English reading proficiency. Participants received \$10 as remuneration, with the top 25% receiving an additional \$5 bonus. The experiment took approximately 25 minutes in total.
6.2. Apparatus
[…]
Boxplot comparison of Manual and Automatic approaches across five measures: Duration, Psychological Ownership, Performance, Effort, and Frequency of Use. Blue data points overlay black boxplots, indicating data distribution and variance.
6.3. Results
6.3.1. Duration
6.3.2. Psychological Ownership
6.3.3. Task Workload
6.3.4. Preferences
Majority (16, 53%) appreciated manual for the increased control it enabled, for example: “I felt a little more control [when] I was able to pick what I wanted to be summarized. [When] I just pressed one button, it summarized everything, but not necessarily the areas I wanted it to”(P22). A few (4, 13%) mentioned how manual provided more incentive to read the document, for example: “generating all summaries at once will encourage users not to actually read the document for themselves. Generating individual summaries after I selected text meant I [had] to actually read the document”(P10).
Overall, our results suggested that selecting text for an AI margin note is slower and requires more effort when done manually. However, this may be a worthwhile trade-off, as manually selecting text was associated with higher feelings of psychological ownership, better perceived performance, and was preferred by participants. In this experiment, the resulting AI margin note was always a summary, however, there are many other types of AI margin notes that can be created that involve the human and AI in different ways.
7. AI Margin Note Techniques
[…]
A series of six comments about the parts of a flower. Each shows different activities: summary, fill-in-the-blank, prompt writing, answering a practice question, feedback on text, and writing without AI. The comments focus on identifying and explaining the parts and functions of flowers: sepals, petals, male parts, and female parts.
To fully explore the design parameter of human and AI involvement, we implemented six techniques that fall along a continuum, ranging from AI margin notes whose text was generated entirely by the LLM, to those whose text was written entirely by the user (Figure 9). All are created by selecting text manually and pressing a blue “Comment” button.
7.1. Receive Summary
The technique with the least human involvement is receiving a generated summary where 100% of the text is written by AI (Figure 9a). This is identical to the manual text selection technique presented in Experiment 2.
7.2. Complete Fill-in-the-Blank Exercise
A technique with slightly more human involvement is receiving a generated summary, but with some keywords omitted to resemble a fill-in-the-blank question (Figure 9b). Like the previous technique, the LLM produces a summary for the text selected by the user, shown in black text and prepended by a black ‘sparkle’ icon. Each summary contains one or two ‘blanks,’ displayed using dropdowns containing three options. The user must select the correct answer from the provided options. Underneath, the comment displays feedback to notify the user whether their responses are correct (shown in green with a check mark) or incorrect (shown in red with an ‘X’). The user can delete the comment or regenerate it.
7.3. Write Prompt
A technique with some human and some AI involvement is writing a custom prompt for the LLM to respond to (Figure 9c). Writing a prompt allows the user to exert more control over the output, for example, to achieve goals that go beyond summarization such as learning additional information about the text or simplifying it. This is identical to the AI margin note technique described in Experiment 1.
7.4. Answer Practice Question
Another technique with some involvement from both the human and the AI is answering a practice reading comprehension question that was generated by the LLM, which has been shown to help with learning (Wang et al., 2024; Liu et al., 2023). The LLM produces a practice short-answer question based on the selected text (Figure 9d), prepended with a grey ‘sparkle’ icon and shown using grey italicized text with a dashed border. The user types in their answer in a text box underneath and presses a blue “Confirm” button to submit their response. The submitted response is then shown in black text. After a few seconds, the user receives feedback from the LLM about their answer: correct responses are shown in green and with a check mark underneath the response, and incorrect responses are shown in red with an ‘X.’ The user can edit their responses to correct mistakes (and will receive new feedback accordingly), or can delete their comment.
7.5. Receive Feedback on Written Text
A technique that has a lot of human involvement is requiring users to write their own comment text, with LLM-generated feedback and suggestions on how their comment can be improved (Figure 9e). This can be beneficial, as many learners struggle to take effective notes (Bretzing and Kulhavy, 1979). The user can type within the provided text box and press “Confirm.” The comment text is shown as black text. After a few seconds, the AI assistant provides feedback on the text, prepended with a grey ‘sparkle’ icon and shown as grey, italicized text with a dashed border. A green check mark is shown if their comment text is good, and a red ‘X’ is shown if it is not. The user can edit their comment to receive updated feedback and can delete it.
7.6. Write Text without AI
The technique with the most human involvement is requiring the user to write their own comment, without any AI assistance (Figure 9f). This baseline resembles the behaviour of existing commenting systems, where 100% of the text is written by the user. The user can type within the provided text box and press “Confirm.” The comment is shown as black text. The user can edit or delete the comment.
[…]
Boxplot comparison of None, Feedback, Question, Prompt, Blank, and Summary across five measures: Duration, Psychological Ownership, Mental Demand, Performance, and Effort. Coloured data points overlay black boxplots, indicating data distribution and variance.
8. Experiment 3: Human and AI Involvement
The goal of this experiment is to investigate the effect of human and AI involvement. Participants read six documents, one for each AI margin note technique described above: receiving a summary (summary), completing a fill-in-the-blank exercise (blank), writing a prompt (prompt), answering a practice question (question), receiving feedback on written text (feedback), and writing without any AI assistance (none). The experiment was longer, approximately 90 minutes.
8.1. Participants
We recruited 36 new participants on Prolific, using the same inclusion criteria as the previous experiments. Four (11%) were excluded for experiencing technical issues, leaving 32 valid responses (Table A.5). All but one were proficient at reading in English. Participants received \$25 total, with an additional \$5 bonus if their total reading comprehension score was within the top 25%.
8.2. Results
Where applicable, we use Friedman omnibus tests and Wilcoxon signed-rank post hoc tests, with Holm’s corrections for multiple comparisons. Statistical test details are shown in Table A.6.
8.2.1. Reading Comprehension
Contrary to what prior work suggests, techniques with more human involvement were not associated with higher reading comprehension scores. Specifically, the differences between summary ($\textsc{mdn}\!=\!4$, $\textsc{iqr}\!=\!2.25$) and all other conditions (all $\textsc{mdn}\!=\!5$, $\textsc{iqr}\!=\!2$) are marginal ($p=.05$), and not statistically significant.
8.2.2. Duration
8.2.3. Psychological Ownership
8.2.4. Task Workload
Generally, techniques that required more human involvement were more mentally demanding, associated with poorer perceived performance, and more effortful. For Mental Demand, there was a significant effect of condition (Figure 10c) and post hoc tests showed that none ($\textsc{mdn}\!=\!61$, $\textsc{iqr}\!=\!47$) and feedback ($\textsc{mdn}\!=\!70.5$, $\textsc{iqr}\!=\!31$) were associated with higher scores than prompt ($\textsc{mdn}\!=\!40.5$, $\textsc{iqr}\!=\!52.25$), blank ($\textsc{mdn}\!=\!20.5$, $\textsc{iqr}\!=\!30.75$), and summary ($\textsc{mdn}\!=\!11.5$, $\textsc{iqr}\!=\!22.25$).
There was also a significant effect of condition on Performance (Figure 10d), with post hoc tests revealing that participants generally felt like they did better at the task with blank ($\textsc{mdn}\!=\!100$, $\textsc{iqr}\!=\!8.5$) than none ($\textsc{mdn}\!=\!90.5$, $\textsc{iqr}\!=\!23.5$), feedback ($\textsc{mdn}\!=\!95$, $\textsc{iqr}\!=\!15.25$), and summary ($\textsc{mdn}\!=\!100$, $\textsc{iqr}\!=\!20.75$).
For Effort, a significant effect of condition (Figure 10e) and post hoc tests showed that none ($\textsc{mdn}\!=\!62.5$, $\textsc{iqr}\!=\!42.75$) and feedback ($\textsc{mdn}\!=\!68$, $\textsc{iqr}\!=\!32.5$) had the highest scores, and summary ($\textsc{mdn}\!=\!10$, $\textsc{iqr}\!=\!36.75$) had the lowest. We did not observe significant differences between the different techniques for Physical Demand, Temporal Demand, and Frustration.
8.2.5. Preferences
9. Discussion
Overall, our results show that participants preferred AI margin notes over traditional, chat-based interactions. When AI margin notes were created through manual text selections, participants took longer to read the document and this required more effort, but they felt more psychological ownership and felt like they did a better job as a result. These feelings may have contributed to the strong preferences they had for manually-created AI margin notes. In short, our results suggest that AI margin notes were valued for their integration, and are perceived to be especially valuable with manual text selection.
Regarding human and AI involvement, our results suggest that different techniques have their own strengths and weaknesses. For example, techniques that required more human involvement, like receiving feedback about written text, required more mental demand and effort, but were associated with higher feelings of psychological ownership than those with low human involvement, like receiving summaries. However, participants preferred techniques with more AI involvement. One promising technique was receiving generated summaries with fill-in-the-blank exercises, the most preferred technique. It was perceived as not too mentally demanding or effortful, still associated with moderate feelings of psychological ownership, and participants felt like they performed well when creating them.
The implications for design are to adopt the integrated AI margin note approach within document reader software, to support manual text selection, and to provide options of AI margin note technique that vary in human and AI involvement, depending on the user’s preferences and goals.
9.1. Reading Comprehension
Given the literature that suggests that increased cognitive engagement can increase learning, we were surprised to see that more cognitively demanding AI margin note techniques were not associated with higher reading comprehension scores. To confirm this is not due to one specific experimental protocol, we conducted other pilot experiments with different participants and key variations. For example, we ran experiments with 24 hour gaps between the reading and testing stages to ensure that the documents were not simply remembered in short-term memory, but we did not observe any significant differences. Participants self-declared their knowledge of the topics discussed in each document, so we tried omitting participants with higher background knowledge, and found similar results. Another possibility is a ceiling effect because the range of possible scores is low (0-6). To rule this out, we conducted another pilot experiment using documents and questions from Guidroz et al. (Guidroz et al., 2025), [^3] We emailed the authors and received permission to reuse their materials. enabling more granular (0-12) scores. These documents were also more difficult (university graduate level, rather than an 8th grade level). Even with this increased score granularity and more difficult reading task, we found similar results, with no significant differences between techniques. Therefore, we believe that less cognitively engaging techniques with more AI involvement, like receiving generated summaries, may not be as detrimental to reading comprehension as one may assume.
9.2. Limitations
Our results, especially results related to reading comprehension, may not hold after extended long-term use. For example, LLMs may hallucinate and produce incorrect AI margin notes, and people may come to over-rely by studying incorrect notes (Wang and Fan, 2025). Techniques with low human involvement may make students ‘lazy’ over time and impact their ability to take notes in situations where LLMs are unavailable.
Our results may not also apply to different types of documents and longer documents. For example, prior work on text highlighting suggests that readers should be able to distinguish between important and unimportant text (Fowler and Barker, 1974), however, readers may struggle to identify the most important information within a long and complex document about an unfamiliar topic. In these cases, automatically selecting text within the document to create AI margin notes may be preferred.
Some aspects of our experimental design may be lacking in ecological validity. Notably, we chose to keep the number of AI margin notes and chat responses constant across techniques, and participants could only interact with one technique at a time. We made these decisions for increased control, however, such restrictions would not exist in a real system, where users would create as many AI margin notes and in whatever way they wish.
We focused on the specific context of creating AI margin notes as a study aid. However, people leave comments in document readers for many other reasons, for example, to leave editorial notes. Other AI margin note designs may be better suited to these goals.
10. Conclusion and Future Work
We propose and explore the design of “AI margin notes” that leverage the commenting feature of document reader software to provide LLM capabilities in a way that is more integrated into document text. Three experiments evaluated variations from different design parameters: integration, selection automation, and human and AI involvement, and overall, participants valued having integrated AI margin notes and creating them manually. AI margin note techniques that involved the human and AI to different degrees were valued for different reasons, suggesting that document reader software should provide multiple variations to support different user goals and preferences.
Future work should incorporate AI margin notes into existing software, explore other types of AI margin notes, and investigate the effects on factors like reading comprehension over the long term and with different types of documents. Our work adds more evidence that chat-based interfaces are not the only way of interacting with LLMs, and that increased integration with document text is beneficial, especially when they are created manually and the trade-offs of human and AI involvement are considered.
Appendix A Appendix
Table A.1. Demographics for Experiment 1.
@@ Appendix A Appendix
[…]
Demographic information: gender, age, education, experience using document readers, commenting experience, LLM experience, and LLM document summary experience. The demographics cover a diverse range.
Table A.2. Statistical test results for Experiment 1.
Measure $W$ $p$ RBCRBC
Reading Comprehension 80 .81 .06
Duration 148 .50 .16
Psychological Ownership 82.5 .40 .21
Mental Demand 149 .72 -.08
Physical Demand 76 .68 .11
Temporal Demand 74.5 .63 -.13
Performance 70.5 .78 .08
Effort 105.5 .32 .24
Frustration 57 .12 -.40
Frequency of Use 77.5 .07 .44
[…]
Details of statistical tests. Significant tests are denoted with asterisks, and more asterisks indicate smaller p-values.
Table A.3. Demographics for Experiment 2.

### v3

@@ 4. Experimental Method
We conducted three experiments to explore these three design parameters of AI margin notes. Specifically, we were interested in understanding how integration, selection automation, and human and AI involvement affect reading comprehension, duration, psychological ownership, task workload, and user preferences. All experiments used the same experimental method and were conducted using the Prolific crowdsourcing platform, [^1] https://www.prolific.com but with different participants for each experiment. Note our protocol was reviewed and approved by our institution’s Research Ethics Board.
@@ 4.1. Task
The primary experimental task was composed of two stages. First, a reading stage, where participants read non-fiction documents that were approximately 500 words each. For each document, they interacted with an LLM assistant using a different technique that represents a variation within the design parameter under evaluation. Second, a test stage, where participants completed reading comprehension tests about each document. Each test consisted of six multiple choice questions. The documents and questions were developed by Wallace et al. (WallaceReadingPassages) and were designed by a learning and reading specialist to be suitable for an eighth grade reading level. [^2] Note that we slightly modified the wording of some questions, as we noticed that some questions could be successfully answered by looking at the wording of other questions. Our versions of these questions and the associated documents are included in the supplementary materials. At the time of writing this paper, Wallace et al.’s license agreement permits reuse, modification, public display, and redistribution for non-commercial research purposes. This is representative of documents targeted for the general public (nngroupLegibilityReadability).
@@ 4.2. Apparatus
Our experimental system was a custom Node.js and React web application that implemented two types of interfaces, one for each stage: the reading interface, where participants read a document and interacted with an LLM using a specific technique (Figure 5); and the testing interface, where participants answered time-bounded, multiple-choice reading comprehension tests. The supplementary video demonstrates both interfaces.
@@ 4.2.1. Reading Interface
The document was displayed on the left side of the screen. The specific design variation of the experimental condition was displayed beside it (details provided in each individual experiment). All design variations that involved an LLM used GPT 4.1 mini. [^3] The system prompts used to generate responses are included in the supplementary materials. The top of the screen contained a toolbar indicating how many comments or responses they still had to complete and a blue “Finish Reading” button to end the trial.
@@ 4.3. Procedure
Participants received a link to an experimental web application through Prolific. They had to use a desktop or laptop computer, which was strictly enforced through the web application. First, participants read a consent form, which detailed inclusion and exclusion criteria, data handling procedures (anonymized and stored on encrypted hard drives), and remuneration. After providing informed consent, participants entered basic demographic information and read instructions. [^4] The demographics questionnaire and all instructions are included in the supplementary materials. The instructions described the general nature of the reading and testing stages, and details on how to use the specific techniques being tested.
@@
Next, they completed the reading stage. There was no time limit, but to focus on the effect of the different techniques rather than the number of responses, participants had to leave 3 AI margin notes (or receive 3 responses from the AI assistant, depending on the technique). After reading the document, they answered 10 questions about their experience and 1 question about their prior knowledge of the document’s content. They repeated this for the other documents, then answered 2 questions about their overall preferences.
@@
Two hours later, the participant was invited to return for the test stage. Given the lower reading difficulty of the documents, we increased the difficulty of the test by making it closed-book and restricted to 60 seconds. After completing all tests, participants described other study aids they used, such as taking a screenshot of the document or writing notes outside of the reading interface.
Table 1. Experiment 1 demographics.
@@
4.4. Design
All experiments use within-subjects design since the documents were relatively short, and we wanted participants to compare their experiences with each technique. There is one independent variable, condition, which was randomly assigned. For a single condition, one of six documents from Wallace et al. (WallaceReadingPassages) was randomly-assigned. For increased internal validity, a single document was not restricted to a single condition and could be assigned to any condition.
4.5. Measures
Reading comprehension and duration were calculated from logs, and all other metrics were calculated from subjective questionnaires completed after the reading stages. With the exception of rankings, these used a 0-100 interval range. [^5] The supplementary materials contains all questions.
4.5.1. Reading Comprehension
4.5.2. Duration
4.5.3. Psychological Ownership
This was measured by asking two quantitative questions about Personal Ownership and Responsibility (Caspi2011Collaboration). The average of the two create a composite measure, as done in prior work (joshi2025interactiontechniquesencouragelonger; ConstrainedGPT). The internal consistency reliability score, which describes how consistently different items on a questionnaire describe the same underlying concept, was high for all experiments (.80 $\leq\alpha\leq$ .90), suggesting that the composite measure was appropriate to use for data analysis. As described previously, this is an important measure since fostering psychological ownership can improve learner experiences.
4.5.4. Task Workload
We used factors from the NASA-TLX (NASATLX): Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Techniques that are more manual or that require more human involvement may involve more mental demand and effort, which could be beneficial for learning (bjork1994memory; bjork2011making; bjork1994institutional; tyler1979cognitive), however, this may not happen if they are too frustrating (grund2024learning).
4.5.5. Preferences
We asked about Frequency of Use, a question from the system usability scale (SUS) (brooke1996sus) representing how frequently the participant would use the feature if made available to them. After trying all techniques, participants also gave each condition a Ranking, where 1 was the best, 2 was the second best, and so on. Ties were allowed. To establish an Overall Ranking of conditions, we use the Condorcet voting method (young1988condorcet). An overall rank of 1 means that technique ‘defeats’ all others in pairwise comparisons. An overall rank of 2 means that technique ‘defeats’ all others except for the technique ranked first, and so on. This is important to study, as users would likely want to keep using techniques that they like, which is especially useful if it also improves their learning outcomes.
4.5.6. Other Metrics
We triangulate these measures with other data, such as characteristics of participant prompts and the length and location of selected text. These metrics are specific to each experiment, so we introduce them with their results.
5. Experiment 1: Integration
5.1. Participants
We recruited 29 participants through Prolific. Participants were restricted to the United States and Canada, and those who had completed 2,500 previous tasks on the platform with a 99-100% approval rating. Three participants (10%) were removed for using other study aids prior to the test, leaving 26 valid responses (Table 1). All self-reported proficiency in reading in English. Each participant received \$10, plus a \$5 bonus for scoring in the top 25% on the comprehension test, to encourage more conscientious active reading. The experiment took roughly 25 minutes in total.
5.2. Apparatus
5.3. Results
We use Wilcoxon signed-rank tests to investigate the effects of condition on the various measures. To streamline the presentation of results, details of these statistical tests are shown in Table A.1 and all data is included in the supplementary materials.
5.3.1. Preferences
Although there were no differences for the aforementioned measures, most participants preferred prompting with AI margin notes. From the Condorcet voting method, we observed that note received an Overall Ranking of 1, and chat received an Overall Ranking of 2. Notably, the majority of participants (17, 65%) assigned note a Ranking of 1, and chat a Ranking of 2 (16, 62%).
To better understand why participants preferred note over chat, we examined participants’ free-form responses. Several (12, 46%) mentioned how note was easier and more intuitive than chat. They described how they “liked [not] having to move to a separate chat”(P21), how “having the comment embedded right into the document [was] handy and easy to reference along with the actual text”(P5), and how “selecting text [was] more natural when one has a question related to that part of the text”(P14).
5.3.2. Prompt Wording
These responses regarding the ease of referring to specific parts of the document were corroborated by examining how participants worded their prompts. Participants did not have to write prompts that were as specific when they used note, perhaps due to the increased specificity enabled by text selections. Notably, 29 prompts (37%) written with note referred to specific nouns in the document using deictic words (e.g., “this,” “they,” “it,” and “here”). With chat, deictic words were less frequent (17, 22%) and most instances referred to specific nouns from previously-entered prompts (10, 13%). For example, consider P3 and P21, who both asked follow-up questions about materials used to make wagon wheels. P21, who read this document with chat, asked “what were wagon wheels made of before iron?” But P3, who used note, instead selected text that discussed wagon wheel material, and asked “how much of an impact did this have for the world?”
5.4. Summary
Overall, our results suggested that the integration of AI margin notes is highly desirable and preferable over chat-based interfaces as it prevented switching between interfaces; it was easier to reference individual comments later; and it was easier to to ask questions using deictic words to refer to specific nouns described in the selected text. In this experiment, associating text to the AI margin note was done manually by the participant. However, selecting text could also be done automatically by the LLM, which we explore further in the following experiment.
Table 2. Experiment 2 demographics.